OpenStudy (anonymous):

what is the eccentricity of the ellipse (x+7)^2/16+(y-3)^2/4=1

OpenStudy (anonymous):

type in into your calculator joonam

OpenStudy (amistre64):

the e value i believe is c/a

OpenStudy (amistre64):

a^2 is the bigger of the 2 denoms; and b^2 the smaller; then to find c we utilize the pythag thrm b^2 + c^2 = a^2

OpenStudy (amistre64):

c^2 = 16 - 4 c^2 = 12 c = 2sqrt(3) e value = 2sqrt(3)/4 = sqrt(3)/2 if i recall it correctly